OpenStudy (anonymous):

A rabbit runs along a straight and narrow path w a constant speed of 20 m/s. He passes a sleeping tortoise, which immediately starts to chase the rabbit w a constant acceleration of 3.5 x 10^-3 m/s^2. how long does it take the tortoise to catch the rabbit? give you answer in hours.

OpenStudy (anonymous):

using v= u + at u=0 20=3.5*10^(-3)t t = 20/3.5*10^(-3) t=0.158*10^4=1.58*10^4 hours (approx)